# MoIRA: Modular Instruction Routing Architecture for Multi-Task Robotics

### v1

@@ 3.3 Routing Strategies
    Embedding Similarity: We embed both task descriptions and expert meta-descriptions using all-MiniLM-L6-v2 (23M parameters) [21]. Cosine similarity determines the routed expert ID.
@@
    Prompt-Driven Routing: SmolLM2-1.7B-Instruct [1] receives structured prompts [^1] SmolLM2 receives a prompt containing meta-descriptions of each expert and a short chain-of-thought instruction set. Example tasks with matched specialists are provided to guide classification. The model is then asked to select the most appropriate expert (e.g., Output: 0, 1, or 2) based on task semantics. containing the task and candidate expert descriptions. The model selects the most likely expert via language model inference.
@@ 3.4 Classification and Execution
We assign each task a corresponding ground-truth category: embodiment type (for GR1) or semantic type (for LIBERO). We formulate routing as a classification problem. Routing accuracy is measured using the macro-averaged F1 score between predicted and ground-truth labels. At inference time:
  • The router selects the appropriate expert based on the input task description.
  • If using dynamic adapter loading, the corresponding LoRA weights are loaded (swap latency: 9.4s).
  • The selected expert then executes the assigned task.
Swap latency can be amortized over multiple episodes to reduce its impact during extended deployments. MoIRA serving and inference configurations are depicted in Figure 3
@@ 3.5 Evaluation Protocol
GR1: We report mean squared error (MSE) of the active joints against the expert observations over 50 evaluation trajectories per task. The main joints are left hand and right hand for full upper body and arms-only, and right hand only for arms & waist embodiment, respectively.
LIBERO: We report Success Rate (SR) over 100 rollouts for subtask category (Spatial and Goal).
We conduct all our experiments on a single NVIDIA RTX A6000 GPU (48GB VRAM). For all training runs, we keep the default hyperparameters and report evaluation results on 5 seeds.
@@ 4.1 GR1 Results
Table 1: GR1 Specialists vs. Baseline comparison (MSE ± std over 50 Trajectories)
@@
We fine-tune each embodiment-specific specialist for 5K steps and compare it with the pretrained GR00t-N1 baseline in Table 1. We observe significant reductions in MSE: Pouring (1.02 $\rightarrow$ 0.008, 123× reduction), CanSort (0.70 $\rightarrow$ 0.010, 70× reduction), and PlacematToPlate (1.67 $\rightarrow$ 0.31, 5× reduction). These results highlight the effectiveness of embodiment-specific specialization.
Table 2: Various GR1 cross-task MSE ± std evaluation scenarios using two routing types and simple-abstract meta-descriptions
@@ 4.1.2 Cross-Task Generalization
Table 2 presents all GR1 task evaluations across specialists and the jointly-tuned generalist. While specialists retain near-optimal performance on their respective tasks, the generalist demonstrates broader coverage but lower precision. For example, 1.187 MSE on PlacematToPlate versus 0.311 for the matching specialist.
Figure 6: GR1 Pouring and CanSort perturbed task description MSE with MiniLM router.
Figure 7: GR1 Pouring and CanSort perturbed task description MSE with SmolLM router.
@@ 4.1.3 Robustness to Instruction Perturbations
To evaluate routing robustness, we semantically perturb [^2] Perturbed instructions are semantically equivalent rephrasings generated manually for each task, varying lexical choices and syntax (e.g., “pick up the pear on the placemat and set it on the plate” vs “transfer the pear from the mat to the plate”). task instructions by rephrasing them with synonymous verbs and alternate phrasing. For example, ”take the bell pepper from the placemat and move it to the plate” may be rewritten as ”grab the bell pepper off the placemat and put it onto the plate”.
Table 3: Router F1 Score and Held-out GR1 Tasks Evaluation (MSE $\pm$ std over 10 episodes per task)
Model WineToCabinet Cuttingboard Coffee F1 (avg)
MiniLM1 1.669 ± 0.047 0.399 ± 0.287 0.201 ± 0.023 1.00
MiniLM2 1.662 ± 0.046 0.294 ± 0.209 0.196 ± 0.018 1.00
SmolLM21 1.669 ± 0.046 0.735 ± 0.677 0.192 ± 0.019 0.87
SmolLM22 1.664 ± 0.047 0.482 ± 0.413 0.197 ± 0.016 0.97
CuttingboardToTieredBasket.
1 abstract meta descriptions; 2 simple meta descriptions
As shown in Table 4 and Figures 6-7, SmolLM2 maintains stable performance across tasks regardless of prompt type. In contrast, MiniLM (abstract) shows sharp degradation (e.g. MSE=0.16 on CanSort), indicating higher sensitivity to linguistic variation.
4.1.4 Generalization to Held-Out Tasks
We evaluate MoIRA on three held-out GR1 tasks. As shown in Table 3, MiniLM (simple) achieves perfect routing (F1 = 1.00) with low MSE across tasks. In contrast, SmolLM2 (abstract) yields lower MSE on CuttingboardToTieredBasket (0.735) but suffers from reduced routing reliability (F1 = 0.87). MoIRA significantly outperforms a jointly-tuned generalist model on two out of three tasks, achieving lower MSE on WineToCabinet (1.55 vs. 1.87) and CuttingboardToTieredBasket (0.41 vs. 1.29), as shown in Table 5. These results support the effectiveness of modular specialization when paired with accurate routing.
Table 4: Router MSE on Original vs. Perturbed Task Descriptions (10 Episodes per Task)
@@ 4.1.5 Routing Accuracy and Control Performance
Accurate routing is essential for effective control. When the routing F1 score is near 1.0, all models yield low MSE. However, control quality deteriorates once routing F1 drops below $\sim$0.9. This trend is most evident in MiniLM (abstract), which frequently misroutes tasks under perturbation. In contrast, SmolLM2, particularly with simple prompts, maintains strong routing fidelity and downstream control performance due to more reliable classification.
Table 5: MoIRA vs. Jointly-Tuned Generalist Performance on Held-out GR1 Tasks (MSE $\pm$ std over 10 episodes per task)
Model Wine Cuttingboard Coffee
MoIRA 1.55 $\pm$ 0.04 0.41 $\pm$ 0.26 0.19 $\pm$ 0.02
Joint Generalist 1.87 $\pm$ 0.05 1.29 $\pm$ 0.06 0.16 $\pm$ 0.01
@@ 4.2 LIBERO Results
4.2.1 Specialist vs Generalist Success Rates
Figure 8 and Table 7 report Success Rates (SR) on the LIBERO Spatial and Goal subsets. The baseline $\pi_{0}$ model without task-specific fine-tuning achieves 0% SR, as it lacks the necessary observation normalization statistics required for evaluation on these tasks.
Table 6: Comparison of MoE Flexibility and Success Rates (SR) on the LIBERO Benchmark
Method Spatial Goal Routing
Tra-MoE + mask 73 78 Internal
MoDE 90 97 Internal
MoIRA (ours) 94 93 External
In contrast, specialists reach 94% (Spatial) and 93% (Goal), while the jointly-tuned generalist achieves 95% and 90%, respectively. These results highlight the importance of correct routing under MoE settings for performing on par or better than a generalist model.
4.2.2 Comparison to Prior MoE Systems
Table 6 compares MoIRA to Tra-MoE and MoDE. MoIRA’s external routing achieves competitive SRs while retaining modularity and expert flexibility. In contrast to internal routing (e.g., token-level gating in MoDE), MoIRA’s design enables episodic routing with adaptable backbones.
Figure 8: Success Rate (SR) comparison on LIBERO subtasks using the $\pi_{0}$-base model. We evaluate a fully finetuned generalist (gray), jointly-trained LoRA expert on spatial and goal tasks (green), and individual LoRA specialists for each subtask (blue and yellow). All models are trained for 30k steps with default hyperparameters.
Reuss et al. [22] achieve 90% SR on LIBERO Spatial and 97% LIBERO Goal using a sparse expert diffusion transformer. In contrast, MoIRA achieves 93.5% weighted-average SR across Spatial and Goal subtasks using modular $\pi_{0}$ specialists routed via a small language model. Unlike MoDE’s internal token-level routing, MoIRA supports episodic routing via fast adapters, making it suitable for deployment-limited settings.
@@ 4.2.2 Comparison to Prior MoE Systems
We also compare our approach with Tra-MoE [27], which achieves 73% SR on LIBERO-Spatial and 78% on LIBERO-Goal using a transformer-based policy with internal MoE blocks. MoIRA outperforms Tra-MoE on both suites while offering broader routing flexibility and simpler adapter management. These results underscore MoIRA’s strength in maintaining task specialization while scaling across domains, outperforming both internal MoE and diffusion-based approaches in success rate and deployment modularity.
4.2.3 Inference Efficiency and Adapter Management
MoIRA performs runtime specialist switching by loading lightweight LoRA adapters. When all adapters are preloaded in GPU memory, switching between experts is effectively instantaneous, but GPU memory usage scales linearly with the number of specialists. Alternatively, dynamic loading of adapters from CPU reduces memory demands to a single backbone and one active adapter but introduces a latency of approximately 9 seconds per swap (for VLA backbones).
The dynamic setup is especially suitable for amortized, asynchronous use cases, such as domestic service robots completing batches of spatial tasks before switching to goal-directed behaviors.
@@ 4.2.4 Routing Fidelity and System-Level Gains
Consistent with GR1 findings, routing accuracy in LIBERO strongly correlates with downstream task success. SmolLM2 achieves perfect classification performance (F1=1.0) in a zero-shot manner, enabling specialist experts to reliably outperform the jointly-tuned generalist. These results underscore that routing fidelity is the critical point for realizing the full potential of modular expert systems.
@@ 5 Discussion
This work presents MoIRA, a modular MoE routing framework for robotic manipulation tasks. It dynamically assigns pretrained VLA specialists based on robot embodiment or task semantics. To our knowledge, it is the first framework to demonstrate benchmark-validated dynamic specialization of VLA models along these axes. MoIRA uses LoRA adapters and natural language routing to enable strong end-to-end control without requiring MoE retraining. Our results show that accurate routing, rather than further finetuning, is key for the performance gains observed with MoIRA. Evaluation across GR1 and LIBERO benchmarks confirms that task-specific experts, when routed correctly.
The system offers several practical benefits. By finetuning specialists on narrow task distributions, MoIRA improves downstream performance while maintaining modularity. Since experts are decoupled, they can be updated independently, enabling incremental development and domain-specific adaptation. Instruction-following models such as SmolLM2 provide reliable zero-shot routing even under perturbed natural language inputs. MoIRA’s modularity enables per-expert optimization, such as distillation or quantization, without modifying the overall architecture. This supports efficient deployment on low-resource agents while preserving specialization and routing flexibility.
Still, MoIRA has several limitations. Dynamic adapter swapping introduces a latency of $\sim$9 seconds, which limits responsiveness in interactive settings. Preloading all adapters into GPU memory mitigates this at the cost of linearly increasing VRAM usage. All evaluations are currently in simulation. Real-world deployment would introduce additional variability through actuation noise, latency, and sensory drift, which may challenge routing stability. Model performance also depends on router accuracy: while SmolLM2 is robust, MiniLM is sensitive to phrasing and ambiguity. The system currently lacks fallback logic; misclassifications cannot be corrected by reverting to a generalist. Additionally, expert meta-descriptions are handcrafted, which creates manual overhead and limits scalability across new task sets.
Another core limitation is MoIRA’s reliance on textual task descriptions for routing. While this design aligns with user-issued commands, it assumes well-formed instructions and ignores contextual signals such as visual observations or environment state. Future work could explore multimodal routing approaches that incorporate perception, history, or task grounding to improve robustness. Real-world validation under physical constraints is a key direction. Incorporating confidence-aware fallback routing could mitigate failures caused by misclassification. Techniques such as adapter caching, speculative routing, and concurrent model loading may reduce latency and improve responsiveness. Another promising avenue is reusing the pretrained language encoder from the VLA backbone as the router, which could eliminate the need for separate models and streamline execution. Finally, while MoIRA relies on manually written meta-descriptions, automating this step using language models could unlock fully autonomous deployment. Generating expert summaries from raw task specs, sensor data, or demonstrations would remove manual bottlenecks and make instruction routing fully end-to-end.
@@
In summary, MoIRA offers a modular, scalable approach to deploying robotic agents that combine embodiment-specific specialization with semantic task understanding. It demonstrates that pretrained experts can be orchestrated through zero-shot lightweight text-description-based routing, enabling generalization without relying on joint training or monolithic models.

### v2

@@ 3.3 Routing Strategies
    Embedding Similarity: We embed both task descriptions and expert meta-descriptions using all-MiniLM-L6-v2 (23M parameters) [20]. The highest cosine similarity among all pairs determines the routed expert ID. This approach treats routing as a geometric retrieval problem, assuming that task instructions and expert capabilities share lexical overlap in a shared latent space. It is computationally negligible but relies on surface-level semantic alignment.
@@
    Prompt-Driven Routing: SmolLM2-1.7B-Instruct [1] receives structured prompts [^1] SmolLM2 receives a prompt containing meta-descriptions of each expert and a short chain-of-thought instruction set. Example tasks with matched specialists are provided to guide classification. The model is then asked to select the most appropriate expert (e.g., Output: 0, 1, or 2) based on task semantics. containing the task and candidate expert descriptions to select the expert via causal language modeling. Unlike embeddings, this leverages the probabilistic reasoning of a LM to approximate $\texttt{P}(\texttt{Expert}_{\texttt{i}}|\texttt{Task},\texttt{Context})$. This allows the router to handle abstract instructions or disparate phrasing where geometric distance may fail, trading inference latency for higher zero-shot routing fidelity.
@@ 3.4 Classification and Execution
We assign each task a corresponding ground-truth category: embodiment type (for GR1) or semantic type (for LIBERO). We formulate routing as a multi-class classification problem, where the router predicts the most appropriate specialist given a natural-language task description. Routing accuracy is measured using the macro-averaged F1 score between predicted and ground-truth labels.
At inference time, MoIRA operates in two sequential stages: routing and expert serving. Given an input task instruction, the router selects an expert index using either embedding-based similarity or prompt-driven language model inference. The selected expert is then served under one of several adapter serving configurations, each exposing a different trade-off between memory footprint and switching overhead.
  • 1.
    Fully instantiated serving. All specialist adapters are kept resident in GPU memory, enabling immediate expert selection. While this configuration minimizes switching overhead, its memory usage scales linearly with the number of specialists, limiting scalability.
  • 2.
    Disk-based swapping. Only a single active adapter is loaded at a time, with inactive specialists stored on CPU or disk and loaded on demand. This approach minimizes GPU memory usage and supports large expert libraries, but incurs significant I/O overhead during expert transitions, making it most suitable for episodic or amortized execution.
  • 3.
    Multi-LoRA serving. An intermediate configuration that maintains multiple lightweight adapters concurrently on a shared backbone, enabling fast expert switching without full model replication. Recent multi-adapter serving frameworks (e.g., S-LoRA [23], LoRAX [19]) exemplify this operating point, balancing memory efficiency with low-latency expert activation.
These serving strategies allow MoIRA deployments to be adapted to hardware constraints and task dynamics. Disk-based swapping enables scalable, storage-bound expert libraries, while multi-LoRA serving expands the operational envelope toward more reactive, low-latency routing. Crucially, MoIRA preserves explicit expert boundaries and supports heterogeneous, independently trained specialists, in contrast to parameter merging or internally gated MoE architectures. MoIRA serving and inference configurations are summarized in Figure 3.
@@ 3.5 Evaluation Protocol
For GR1, we report mean squared error (MSE) of the active joints against the expert observations over 50 evaluation trajectories per task. The main joints are left hand and right hand for full upper body and arms-only, and right hand only for arms & waist embodiment, respectively.
For LIBERO, we report Success Rate (SR) over 100 rollouts for subtask category (Spatial and Goal). We conduct all our experiments on a single NVIDIA RTX A6000 GPU (48GB VRAM). For all training runs, we keep the default hyperparameters and report evaluation results on 5 seeds.
To ensure robust comparisons, we assess statistical significance using Welch’s t-test for continuous metrics (GR1 MSE) to account for unequal variances between specialists and generalists. For binary success rates (LIBERO), we employ Fisher’s Exact Test. We adopt a standard significance threshold of $p<0.05$ for all claims.
@@ 4.1 GR1 Results
Table 1: GR1 Specialists vs. Baseline comparison (MSE ± std over 50 Trajectories).
@@
We fine-tune each embodiment-specific specialist for 5K steps and compare it with the pretrained GR00t-N1 baseline in Table 1. We observe significant reductions in MSE: Pouring (1.02 $\rightarrow$ 0.008, 123× reduction), CanSort (0.70 $\rightarrow$ 0.010, 70× reduction), and PlacematToPlate (1.67 $\rightarrow$ 0.31, 5× reduction). We find the reduction in error to be statistically significant across all tasks ($p<0.001$ for all 3 tasks), confirming that the reported performance gains are robust to rollout variance and not artifactual.
Table 2: Various GR1 cross-task MSE ± std evaluation scenarios using two routing types and simple-abstract meta-descriptions.
@@ 4.1.2 Cross-Task Generalization
Table 2 presents all GR1 task evaluations across specialists and the jointly-tuned generalist. The specialists demonstrated statistically significant reductions in MSE across all evaluated tasks. Specifically, the improvement was significant for Pouring ($p<0.01$) and highly significant for CanSort and PlacematToPlate ($p<0.001$). Modular specialists achieve superior precision that is statistically distinguishable from the generalist baseline.
Table 3: MoIRA vs. Jointly-Tuned Generalist Performance on Held-out GR1 Tasks (MSE $\pm$ std over 10 episodes per task).
Model Wine Cuttingboard Coffee
MoIRA 1.55 $\pm$ 0.04 0.41 $\pm$ 0.26 0.19 $\pm$ 0.02
Joint Generalist 1.87 $\pm$ 0.05 1.29 $\pm$ 0.06 0.16 $\pm$ 0.01
CuttingboardToTieredBasket.
@@ 4.1.3 Robustness to Instruction Perturbations
To evaluate routing robustness, we semantically perturb [^2] Perturbed instructions are semantically equivalent rephrasings generated manually for each task, varying lexical choices and syntax (e.g., “pick up the pear on the placemat and set it on the plate” vs. “transfer the pear from the mat to the plate”). task instructions by rephrasing them with synonymous verbs and alternate phrasing. For example, "take the bell pepper from the placemat and move it to the plate" may be rewritten as "grab the pepper off the placemat and put it onto the plate".
As shown in Table 4 and Figure 7, SmolLM2 maintains stable performance across tasks regardless of prompt type. In contrast, MiniLM (abstract) shows sharp degradation (e.g. MSE=0.16 on CanSort, Figure 6 ), indicating higher sensitivity to linguistic variation.
Table 4: Router MSE on Original vs. Perturbed Task Descriptions (10 Episodes per Task).
@@
4.1.4 Generalization to Held-Out Tasks
We evaluate MoIRA on three held-out GR1 tasks. MoIRA significantly outperforms a jointly-tuned generalist model on two out of three tasks, achieving lower MSE on WineToCabinet (1.55 against 1.87) and CuttingboardToTieredBasket (0.41 against 1.29), as shown in Table 3. On held-out tasks, MoIRA significantly outperformed the generalist on WineToCabinet and CuttingboardToTieredBasket ($p<0.001$), demonstrating robust zero-shot routing to unseen scenarios. For the Coffee task, the generalist maintained a slight statistical edge ($p<0.01$), though the absolute difference in MSE was marginal ($0.03$).
@@ 4.1.5 Routing Accuracy and Control Performance
Accurate routing is essential for effective control. As shown in Table 5, MiniLM (simple) achieves perfect routing (F1 = 1.0) with low MSE across tasks. In contrast, SmolLM2 (abstract) yields lower MSE on CuttingboardToTieredBasket (0.735) but suffers from reduced routing reliability (F1 = 0.87). When the routing F1 score is near 1.0, all models yield low MSE. However, control quality deteriorates once routing F1 drops below $\sim$0.9. This trend is most evident in MiniLM (abstract), which frequently misroutes tasks under perturbation. In contrast, SmolLM2, particularly with simple prompts, maintains strong routing fidelity and downstream control performance due to more reliable classification.
Table 5: Router F1 Score and Held-out GR1 Tasks Evaluation (MSE $\pm$ std over 10 episodes per task).
Model WineToCabinet Cuttingboard Coffee F1 (avg)
MiniLMa 1.662 ± 0.046 0.294 ± 0.209 0.196 ± 0.018 1.00
MiniLMb 1.669 ± 0.047 0.399 ± 0.287 0.201 ± 0.023 1.00
SmolLM2a 1.664 ± 0.047 0.482 ± 0.413 0.197 ± 0.016 0.97
SmolLM2b 1.669 ± 0.046 0.735 ± 0.677 0.192 ± 0.019 0.87
@@
a simple meta descriptions; b abstract meta descriptions
Figure 6: MiniLM Routing Sensitivity. MSE results on GR1 tasks using embedding-based routing. The plot compares performance using simple versus abstract meta-descriptions. While effective on original instructions, the router’s performance shows possible degradation on perturbed instructions, particularly when relying on abstract descriptions. This indicates sensitivity to phrasing variations and potential limitations in semantic robustness.
Figure 7: SmolLM2 Routing Sensitivity. MSE results using the prompt-driven language model router. In contrast to the embedding baseline, SmolLM2 maintains low error rates across both original and perturbed instructions for both tasks. This suggests stronger zero-shot reasoning capabilities and stability against linguistic perturbations.
@@ 4.2 LIBERO Results
Figure 8 and Table 6 report success rates on LIBERO Spatial and Goal subsets. First, we note that the zero-shot pre-trained $\pi_{0}$ baseline yields 0% SR, as it lacks the requisite observation normalization statistics to navigate through episodes successfully. To contextualize the performance of MoIRA’s specialists, we compare three distinct fine-tuning regimes, trained for 30k steps each:
  1. 1.
    Global Generalist (Figure 8, gray bar): a monolithic baseline fine-tuned on the complete LIBERO dataset (gray bars). This represents the performance upper bound of a strong generalist.
  2. 2.
    Joint LoRA Generalist (Figure 8, green bar): a single adapter trained jointly on both Spatial and Goal tasks (green bars), representing a parameter-efficient multi-task baseline.
  3. 3.
    Decoupled Specialists (Figure 8, blue and yellows bars): independent LoRA adapters fine-tuned exclusively on their respective subsets (blue and yellow bars), routed via MoIRA.
@@
4.2.1 Specialist vs Generalist Success Rates
While the jointly-tuned generalist achieves marginally higher raw success rates (95% vs 94% Spatial; 90% vs 93% Goal), there was no statistically significant difference revealed between the models ($p>0.05$). Specifically, for Spatial tasks ($p=1.0$) and Goal tasks ($p\approx 0.61$), the 95% confidence intervals significantly overlap. These results statistically confirm performance parity – MoIRA matches the generalist’s SR while offering the architectural benefits of modularity and reduced inference memory.
Table 7: Comparison of MoE Flexibility and SR on LIBERO Benchmark. MoIRA outperforms Tra-MoE and achieves parity with MoDE.
Method Spatial Goal Routing
Tra-MoE + mask 73 78 Internal
MoDE 90 97 Internal
MoIRA (ours) 94 93 External
Individual specialists reach 94% Spatial and 93% Goal success rate, while the jointly-tuned generalist achieves 95% and 90%, respectively. Precise MoE routing is an important aspect of retaining parity with a generalist model.
4.2.2 Comparison to Prior MoE Systems
Table 7 compares MoIRA to Tra-MoE [28] and MoDE [21]. MoIRA’s external routing achieves competitive SR while retaining modularity and expert flexibility. In contrast to internal routing (e.g., token-level gating in MoDE), MoIRA’s design enables episodic routing with adaptable backbones.
Figure 8: Comparative Success Rates (SR) on LIBERO subtasks. We contrast a Global Generalist (gray; full model fine-tuning on the entire dataset) against a Joint LoRA Generalist (green; single adapter trained on combined Spatial and Goal data) and MoIRA-routed Specialists (blue/yellow; decoupled adapters for each subtask). The modular specialists achieve competitive performance with the monolithic baselines despite being trained on strictly partitioned data.
Reuss et al. [21] achieve 90% SR on LIBERO Spatial and 97% LIBERO Goal using a sparse expert diffusion transformer. MoIRA scores 93.5% weighted-average SR across Spatial and Goal subtasks and achieves statistical parity with MoDE (93.5% vs 93.5%), matching state-of-the-art diffusion policies using a modular VLA architecture. Unlike MoDE’s internal token-level routing, MoIRA supports episodic routing via fast adapters, making it suitable for deployment-limited settings.
We also compare our approach with Tra-MoE [28], which achieves 73% SR on Spatial and 78% SR on Goal using a transformer-based policy with internal MoE blocks. MoIRA outperforms Tra-MoE on both suites while offering broader routing flexibility and simpler adapter management.
@@ 4.2.3 Routing Fidelity and System-Level Gains
Consistent with GR1 findings, routing accuracy in LIBERO strongly correlates with downstream task success. SmolLM2 achieves perfect classification performance (F1=1.0) in a zero-shot manner, enabling specialist experts to reliably outperform the jointly-tuned generalist. This underscores that routing fidelity is the critical point for realizing the full potential of modular expert systems.
@@ 5 Discussion
This work presents MoIRA, a modular MoE routing framework that dynamically assigns pretrained VLA specialists based on robot embodiment or task semantics. To our knowledge, it is the first framework to demonstrate benchmark-validated dynamic specialization of VLA models along these axes. MoIRA uses LoRA adapters and natural language routing to enable strong end-to-end control without requiring MoE retraining. Our results demonstrate that this modular architecture achieves performance parity with monolithic generalists and other robotic MoE systems while significantly enhancing scalability and deployment efficiency. By decoupling expert specialization from orchestration, MoIRA validates that independent, lightweight adapters offer a flexible and resource-efficient alternative to rigid joint-training pipelines.
Sim-to-Real and Sensory Robustness. Current evaluations are performed in simulation. While this allowed us to validate algorithmic stability across highly distinct morphologies (GR1) and semantic domains (LIBERO), real-world deployment introduces actuation noise and sensory drift. The router’s reliance on purely textual descriptions currently assumes well-formed instructions and ignores environmental context (e.g., distinguishing between two identical cups). To address this, future work will explore multimodal routing by integrating lightweight vision encoders (e.g., CLIP) or extracting features from the frozen VLA backbone itself. This would allow the router to disambiguate tasks based on visual state without sacrificing the architecture-agnostic nature of the external controller.
Latency and Operational Envelope. A key trade-off in our design is the I/O latency introduced by standard dynamic adapter loading. This characteristic typically restricts MoIRA to episodic or batch-processing workflows, such as a service robot completing a 20-minute ”kitchen cleaning” sequence before switching to ”laundry folding”, rather than high-frequency, reactive switching. To address this bottleneck, we integrated specialized multi-LoRA serving frameworks, which facilitate concurrent serving of fine-tuned adapters with millisecond-level switching overhead. While this approach effectively masks latency and expands the operational envelope for a subset of compatible backbones, further investigation into universal architecture support and speculative routing strategies remains a priority for fully reactive deployment.
Scalability and Automation. Finally, while SmolLM2 demonstrates robust zero-shot routing, the reliance on manually written meta-descriptions limits rapid scaling to thousands of experts. Automating this pipeline is a high-priority research direction. We envision using large language models to auto-generate expert summaries from raw task specifications or demonstration logs, effectively closing the loop to create a fully autonomous, self-organizing mixture of robotic experts.
@@
In summary, MoIRA offers a modular, scalable approach to deploying robotic agents that combine embodiment-specific specialization with semantic task understanding. It demonstrates that pretrained experts can be orchestrated through zero-shot lightweight text-based routing, enabling generalization without relying on joint training or monolithic models.